You mix a 105.0 −mL sample of a solution that is 0.0130 M in NiCl2 with a 185.0 −mL sample of a solution that is 0.300 M in NH3.

After the solution reaches equilibrium, what concentration of Ni2+(aq) remains? The value of Kf for Ni(NH3)62+ is 2.0×108.
Answer #1

Ni2+ + 6NH3 ---> [Ni(NH3)6]2+

mmol of Nickel ion =105 x0.013=1.365

mmol of ammonia =185x0.3=55.5

total vol=290ml

this means 1 mmol of Ni ion react to 6 mmol of ammonia.

so 1.365 mmol will react to=1.365x6=8.19mmol

as ammonia is in excess so 8.19mmol complex is form=8.19/290=0.0282M

[Ni(NH3)6]2+ ---> Ni2+ + 6NH3

means 2x108 = C-x/x .x7 =c/x8 =0.0282/x8

it will give value of x which is Ni ion conc finally